OpenStudy (anonymous):

a video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. on level 1, a player must earn 1 experience point. on level 2, a player must earn 7 experience point. on level 3, a player must earn 15 experience point. on level 4, a player must earn 25 experience point. How many experience points must a player earn on level 7?

OpenStudy (anonymous):

its 67

OpenStudy (anonymous):

The difference: Start with 6: 1 + 6 = 7 then increasing: 7 + 8 = 15 15 +10 = 25 25 + 12 = 37 .....